# Assessing Inf luential Observations in Pain Prediction using fMRI Data

### v1

@@ 4.2 Assessment Metrics
We consider three assessment criteria, which are the power of detection, false positive rate and algorithm running time. Denoted by $\text{Power}_{\text{Detection}(\text{Selector})}$, $\text{FPR}_{\text{Detection}(\text{Selector})}$ and $\text{Time}_{\text{Detection}(\text{Selector})}$, they are defined by
$\text{Power}_{\text{Detection}(\text{Selector})}=\frac{\sum_{b=1}^{1000}\ \widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector}),b}\ \cap\ \{1,...,n_{\text{infl}}\}}{n_{\text{infl}}\times 1000},$
$\text{FPR}_{\text{Detection}(\text{Selector})}=\frac{\sum_{b=1}^{1000}\ \widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector}),b}\ \cap\ \{n_{\text{infl}}+1,...,n\}}{(n-n_{\text{infl}})\times 1000},\ \text{and}$
$\text{Time}_{\text{Detection}(\text{Selector})}=\frac{\sum_{b=1}^{1000}\ \widehat{\text{T}}_{\text{Detection}(\text{Selector}),b}}{1000},$
where “Detection(Selector)” specifies the identification procedure “Detection” when a submodel is selected by “Selector”. Moreover, $\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector}),b}$ and $\widehat{\text{T}}_{\text{Detection}(\text{Selector}),b}$ respectively denote the estimated index set of influential observations and execution time (in seconds) for identifying them based on the $b^{\text{th}}$ generated random sample, $b=1,...,1000$.
4.3 Methods, Codes and Implementation
In our simulations, different combinations of detection procedures and model selectors are considered according to the types of data generating mechanisms.
Under the linear regression model (1), we consider three detection procedures: the ClusMIP (Algorithm 2), and the two competing proposals DF(LASSO) and MIP. In case of $\textbf{X}_{\text{Tp(0.5)}}$ and $\textbf{X}_{\text{Tp(0.8)}}$, we combine the ClusMIP with the LASSO, scaled LASSO (SLASSO), elastic net (ENET), SCAD and MCP, while the ENET is excluded in case of $\textbf{X}_{\text{Tp(0)}}$.
In addition, tuning parameter for all the regularized model selectors is selected via the 10-fold cross-validation (CV), unless otherwise specified. A summary of methods and implementation details is provided in Section 10.2 of the Supplementary Materials.
@@ 4.4 Simulation Results
Now, we present and analyze selected simulation results classified according to the types of data generating mechanisms and perturbation models, for which distinctive patterns of the three assessment benchmarks presented in Section 4.2 have been recognized. All the simulation diagrammatic results are provided in Section 10.3 of the Supplementary Materials.
Linear Regression Model: Perturbation Model I (Figure 2)
Figure 2: Assessment under the linear regression model, Perturbation Model I, $\kappa_{\text{Linear},\textbf{Y}}=$5 and $\textbf{X}_{\text{Tp(0)}}$.
@@
    Power: when $\kappa_{\text{Linear},\textbf{Y}}=5$ and $\textbf{X}=\textbf{X}_{\text{Tp(0)}}$, $\text{Power}_{\text{ClusMIP}(\text{LASSO})}$, $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ and $\text{Power}_{\text{ClusMIP}(\text{SCAD})}$ are higher than $\text{Power}_{\text{MIP}}$ and $\text{Power}_{\text{DF}(\text{LASSO})}$ for the same contamination proportion $\zeta$. Specifically, $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ is the highest, reaching almost 1.00 for all contamination proportions except for $5\%$. In comparison, large variation exists for both $\text{Power}_{\text{MIP}}$ and $\text{Power}_{\text{DF}(\text{LASSO})}$, dropping respectively from 0.83 to 0.32 and 0.64 to 0.19 as $\zeta$ increases. For the remaining simulation configurations, in general we have $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}\approx\text{Power}_{\text{MIP}}\approx 1.00>\big{\{}\text{Power}_{\text{ClusMIP}(\text{LASSO})},\text{Power}_{\text{ClusMIP}(\text{ENET})},\\ \text{Power}_{\text{ClusMIP}(\text{SCAD})},\text{Power}_{\text{ClusMIP}(\text{MCP})}\big{\}}\gg\text{Power}_{\text{DF}(\text{LASSO})}$. Both ClusMIP(SLASSO) and MIP almost capture all the influential points, implying that SLASSO is sensitive to influential points in variable selection, serving as a reasonable candidate for identification purpose. Moreover, detection powers of ClusMIP(LASSO), ClusMIP(ENET), ClusMIP(SCAD) and ClusMIP(MCP) are noticeably lower, ranging between 0.70 and 0.90. In particular, detection powers of ClusMIP(LASSO) and ClusMIP(SCAD) are comparable in magnitude over the same $\zeta$, followed by ClusMIP(ENET) and ClusMIP(MCP). In contrast, $\text{Power}_{\text{DF}(\text{LASSO})}$ is almost always less than 0.4, highlighting its incapability to detect multiple influential observations.
@@
    FPR: across all the design matrices, perturbation scales and contamination proportions, all procedures exhibit consistently small FPR, ranging between 0 and $5.0\%$ which is less than the nominal level $\alpha=0.05$. This phenomenon justifies Theorem 3 for the ClusMIP procedure and Theorem 4 in Zhao et al. (2019) in controlling the FPR. Specifically, the high-dimensional clustering is effective to separate the clean data from the contaminated portion, limiting the possibility for false identification in the subsequent refinement step. On the other hand, the $\text{FPR}_{\text{ClusMIP(SLASSO)}}$ computed for $\textbf{X}_{\text{Tp(0)}}$, $\kappa_{\text{Linear},\textbf{Y}}=5$ and $\zeta=5\%$ seems to be an outlying case, which attains an elevated level of approximately $15.0\%$.
@@
    Computation Time: in general, we have $\text{Time}_{\text{MIP}}<\text{Time}_{\text{DF}(\text{LASSO})}\{\text{Time}_{\text{ClusMIP}(\text{LASSO})},\\ \ll\text{Time}_{\text{ClusMIP}(\text{SLASSO})},\text{Time}_{\text{ClusMIP}(\text{SCAD})},\text{Time}_{\text{ClusMIP}(\text{MCP})}\}\ll\text{Time}_{\text{ClusMIP}(\text{ENET})}$. Noticeably, $\text{Time}_{\text{MIP}}$ and $\text{Time}_{\text{DF}(\text{LASSO})}$ are negligible comparing with that of the ClusMIP approach. This is because that the MIP involves no model fitting. Moreover, the authors further remark [Zhao et al. (2019), Theorem 4] that one iteration of their algorithm suffices to capture all the influential points at the designated false discover rate. Furthermore, the DF(LASSO) requires $(n+1)$ times of model fitting. In contrast, the ClusMIP framework requires $|\widehat{\text{S}}_{\text{infl}}|(2+|n-\widehat{\text{S}}_{\text{infl}}|)$ times of model fitting, where $\widehat{\text{S}}_{\text{infl}}$ is the number of estimated influential points from the high-dimensional clustering scheme. This substantially surpasses that for the DF(LASSO). On the other hand, within the ClusMIP framework, the speed varies according to the forms of penalty functions and particular $\tt{R}$ implementation algorithms. Specifically, the ENET includes a two-dimensional exhaustive search of the tuning parameter pair, leading to the largest computation time. In comparison, the SLASSO simultaneously estimates the noise and coefficient, and its $\tt{R}$ package $\tt{scalreg}$ provides faster convergence than other penalized model selectors under consideration.
Linear Regression Model: Perturbation Model II (Figures 3 and 4)
Figure 3: Assessment under the linear regression model, Perturbation Model II, $\kappa_{\text{Linear},\textbf{X}}$=5 and $\textbf{X}_{\text{Tp(0)}}$.
Figure 4: Assessment under the linear regression model, Perturbation Model II, $\kappa_{\text{Linear},\textbf{X}}$=5 and $\textbf{X}_{\text{Tp(0.5)}}$.
@@
    Power: under all types of design matrices, perturbation scales and contamination proportions, the detection powers of the two existing methods, MIP and DF(LASSO) are no more than 0.11, uniformly and considerably falling behind those of the ClusMIP procedure. This indicates that both approaches may not be able to accurately capture multiple influential points under this specific perturbation formulation. While the under-performance for the DF(LASSO) lies in its leave-one-out design, for MIP, we postulate that augmenting disturbance to the predictors while keeping the response vector intact is unlikely to impose sufficient impact on the marginal correlation-based HIM measure. Thus, true influential points are masked from being detected, leading to low detection power.
@@
    Power: when $\kappa_{\text{Linear},\textbf{X}}$=5 and $\textbf{X}=\textbf{X}_{\text{Tp(0)}}$, we observe that $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$,
[…]
@@
    Power: when $\kappa_{\text{Linear},\textbf{X}}$ increases to 10 and for all types of design matrices, all the $\text{Power}_{\text{ClusMIP}(\text{LASSO})}$, $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$, $\text{Power}_{\text{ClusMIP}(\text{SCAD})}$ and $\text{Power}_{\text{ClusMIP}(\text{MCP})}$ approximately attain 1.00 except for $\text{Power}_{\text{ClusMIP}(\text{LASSO})}$ at $\zeta=15\%$ and $20\%$ for $\textbf{X}_{\text{Tp(0)}}$. In contrast, $\text{Power}_{\text{ClusMIP}(\text{ENET})}$ falls far behind, which is no more than 0.40.
  4. 4.
    FPR and Computation Time: the patterns for FPR and Computation Time are generally analogous to the corresponding ones identified for Perturbation Model I.
Linear Regression Model: Perturbation Model III
Under all the types of design matrices, perturbation scales and contamination proportions considered in our simulations, the generic relationships in detection power, FPR and computation time are similar to the ones identified under Perturbation Model I. In summary, only $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ and $\text{Power}_{\text{MIP}}$ uniformly reach a level between 0.90 and 1.00, whereas $\text{Power}_{\text{DF}(\text{LASSO})}$ is consistently less than 0.10. In comparison, the detection powers of the ClusMIP for the remaining model selectors fluctuate in between 0.20 and 0.80. In addition, as the perturbation scale increases, we observe improved detection powers for all the procedures.
@@ 5 Analysis of Neurologic Signature of Physical Pain
In this section, based on the high-dimensional linear regression model (1), we analyze the performances of the aforementioned influential point detection procedures on the real dataset extracted from Lindquist et al. (2017) on pain level prediction via the fMRI data. Its background and characteristic is provided in Section 5.1. This is followed by presenting the assessment metrics, methods and implementation in Section 5.2. Analysis of results is provided in Section 5.3.
5.1 Background and Characteristic
[…]
The structural T1-weighted images were co-registered to the mean functional image for each subject using the iterative mutual information-based algorithm implemented in the statistical parametric mapping (SPM) software package (Ashburner & Friston, 2005), and were then normalized to the Montreal Neurological Institute (MNI) space. In each functional dataset, initial volumes were removed to allow for stabilization of image intensity. Prior to processing functional images, we removed volumes with signal values that were outliers within the time series. This was done by first computing the mean and the standard deviation of intensity values across each slice, for each image, and then computing the Mahalanobis distances for the matrix of slice-wise mean and standard deviation values. Values with a significant $\chi^{2}$-value (corrected for multiple comparisons based on the more stringent of either false discovery rate or Bonferroni methods) were considered outliers. Next, functional images were slice-time corrected and motion-corrected using SPM. Functional images were warped to SPM’s normative atlas (warping parameters estimated from co-registered, high-resolution structural images) andinterpolated to 2 $\times$ 2 $\times$ 2 $\text{mm}^{3}$ voxels.
5.2 Assessment Metrics, Methods and Implementation
Three assessment metrics are included in our analysis, which are (i) $\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}$, the estimated index set of influential points detected by the procedure “Detection” when a submodel is chosen by “Selector”, (ii) $\big{|}\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector}})\big{\rvert}$, the cardinality of $\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}$, and (iii) $\widehat{\text{T}}_{\text{Detection}(\text{Selector})}$, the time (seconds) required for obtaining $\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}$.
[…]
5.3 Analysis of Real Datasets
The assessment metrics of detection procedures are reported in Table 1, where we make the following observations and analyses:
  1. 1.

### v2

@@ 5.2 Assessment Metrics
Two groups consisting of five metrics are examined. The first group contains the power ($\text{Power}_{\text{Detection}(\text{Selector})}$), false positive rate of detection ($\text{FPR}_{\text{Detection}(\text{Selector})}$), and computation time ($\text{Time}_{\text{Detection}(\text{Selector})}$). Here “Detection(Selector)” specifies the identification procedure “Detection” when a submodel is selected by “Selector”. They directly evaluate the detection performance. The second group includes the MSE of prediction and the empirical probability that the pre-specified true model is contained in the selected submodel (abbreviated as above), each computed both before and after removal of detected influential points for comparison purposes. They highlight the benefits of outlier detection on downstream inference. Here, since MIP has no model selection, the latter two metrics are excluded. These five metrics are formally defined in Section 11.2 of the Supplementary Materials Part I.
5.3 Methods and Implementation
Three detection methods are included: the ClusMIP, DF(LASSO) and MIP. In case of $\textbf{X}_{\text{Tp(0.5)}}$, $\textbf{X}_{\text{Tp(0.8)}}$ and $\textbf{X}_{\text{Sc(0.8)}}$, the ClusMIP is combined with LASSO, scaled LASSO (SLASSO), elastic net (ENET), SCAD and MCP, while the ENET is excluded for $\textbf{X}_{\text{Tp(0)}}$. Moreover, tuning parameter for all the regularized model selectors is chosen via the 10-fold cross-validation (CV), unless otherwise specified. A summary of methods and implementation is given in Section 11.3 of the Supplementary Materials Part I.
@@ 5.4 Simulation Results
Simulation results with distinctive patterns are presented. Others providing similar conclusion are in Sections 1 of the Supplementary Materials Part II for $(n,p)=(100,1000)$ and III for $(n,p)=(160,480)$, respectively.
Perturbation Scheme I (Figure 1)
Figure 1: Perturbation Scheme I with $(n,p)=(100,1000)$,$\kappa_{\text{Linear},\textbf{Y}}=$5 and $\textbf{X}=\textbf{X}_{\text{Sc(0.8)}}$. Panel (A): power of detection; Panel (B): FPR of detection; Panels (C) and (D): MSE in prediction and probability that a selected submodel contains the true model before and after the removal of detected influential observations; Panel (E): algorithm running time in seconds.
@@
    Power: $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ and $\text{Power}_{\text{MIP}}$ are generally higher (above 0.90) than the other procedures for the same contamination proportion $\zeta$ at moderate and high contamination scales $\kappa_{\text{linear},\textbf{Y}}$. This implies that SLASSO is sensitive to influential observations on variable selection, serving as a feasible candidate for use together with an outlier detection procedure. At low values of $\kappa_{\text{linear},\textbf{Y}}$, concerning under-performance was observed under certain scenarios for the MIP. For example, $\text{Power}_{\text{MIP}}$ drops from 0.20 to 0.10 as $\zeta$ increases from $15\%$ to $20\%$; see Panel (A) of Figure 1. In comparison, $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$ at $\zeta=15\%$ and $20\%$ consistently exceeds 0.97. Moreover, within the ClusMIP procedure, the detection powers of ClusMIP applied to LASSO, ENET, SCAD and MCP show higher variation and are generally lower than $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$, ranging between 0.45 and 0.89 under the particular simulation setting shown in Figure 1. In contrast, $\text{Power}_{\text{DF}(\text{LASSO})}$ (¡ 0.38 in Figure 1) is consistently among the lowest due to its leave-one-out design, indicating its limited ability to detect multiple influential points. In summary, ClusMIP(SLASSO) consistently exhibits better detection result than competing procedures, while under-performance of MIP is observed at low contamination intensity. Further, DF(LASSO) appers to be always outperformed in terms of detection.
@@
    FPR: All procedures consistently exhibit small FPR (less than the nominal level $\alpha=0.05$). This phenomenon justifies Theorem 3 of the ClusMIP procedure and Theorem 4 in [39] in controlling the FPR. Specifically, the choices of high-dimensional clustering discussed in Section 4.2.2 are effective in separating the clean and the contaminated portions, limiting the possibility for false detection in subsequent steps.
@@
    MSE of Prediction: Lower prediction MSE after removing detected influential observations is consistently observed across all procedures, contamination proportions, and scales. For example, the MSE of ClusMIP(SLASSO) and DF(LASSO) declines from 35.94 to 2.84 and from 33.23 to 24.18, respectively, at $\zeta=20\%$ as seen in Panel (C) of Figure 1. While the extent of reduction in MSE varies according to the detection procedure, model selector, and contamination scales, MSE is always reduced after eliminating detected outliers, indicating the benefit of their detection. Moreover, low MSE after outlier removal is linked to improved detection performance. Noticeably, the MSE of SLASSO after the removal of outliers is always among the lowest, due to its strong detection performance. In comparison, the MSE after outlier removal for the remaining ClusMIP procedures showshigher fluctuation across simulation settings but generally corresponds to their detection power. For example, in Panels (A) and (C) of Figure 1, the MSE of ClusMIP(SCAD) after outlier removal (between 2.22 and 3.24) is the second lowest which agrees with the second-best position of its detection power performance (between 0.77 and 0.90). In contrast, the MSE of the DF(LASSO) after outlier removal is higher than that of the ClusMIP procedures due to its worse outlier detection.
  4. 4.
    Variable Selection: Improvement in the empirical probability that the true model is included in the selected submodel after removing the detected influential points is generally noted for all the procedures, indicating that these observations are indeed influential on variable selection. Similar to the reduction in MSE discussed above, the performance of this assessment metric hinges on the detection outcome. Due to reduction in the heterogeneity in the data, better detection performance contributes to higher probability of choosing a submodel containing the true model after outlier removal. This is analogous to results seen for detection power. This is numerically corroborated by comparing Panels (A) and (D) of Figure 1: ClusMIP(SLASSO) attains the highest detection power (0.97 – 0.99), followed by ClusMIP(SCAD) (0.77 – 0.90). On the other hand, while the selection probability is approximately 0 for both before outlier removal (i.e., based on the contaminated dataset), the two procedures achieve the highest selection probabilities after outlier removal, ranging between 0.90 – 1.00 and 0.66 – 0.86, respectively. In contrast, the selection probability after outlier removal for DF(LASSO) is 0 at $\zeta=10\%$, $15\%$ and $20\%$, which can be explained by its corresponding poor performance in detection.
  5. 5.
@@
    Computation Time: in general, we have $\text{Time}_{\text{MIP}}<\text{Time}_{\text{DF}(\text{LASSO})}<\{\text{Time}_{\text{ClusMIP}(\text{LASSO})},\\ \ll\text{Time}_{\text{ClusMIP}(\text{SLASSO})},\text{Time}_{\text{ClusMIP}(\text{SCAD})},\text{Time}_{\text{ClusMIP}(\text{MCP})}\}\ll\text{Time}_{\text{ClusMIP}(\text{ENET})}$. Noticeably, $\text{Time}_{\text{MIP}}$ and $\text{Time}_{\text{DF}(\text{LASSO})}$ are negligible comparing with that of the ClusMIP approach. This is due to the fact thatMIP requires no model fitting. Moreover, DF(LASSO) requires models to be fit $(n+1)$ times. In contrast, the ClusMIP framework requires models to be fit $|\widehat{\text{S}}_{\text{infl}}|(2+|n-\widehat{\text{S}}_{\text{infl}}|)$ times, where $\widehat{\text{S}}_{\text{infl}}$ is the number of estimated influential points from the high-dimensional clustering scheme. This substantially surpasses that for the DF(LASSO). On the other hand, within the ClusMIP framework, the speed varies according to the types of penalty functions and particular $\tt{R}$ implementation algorithms. Specifically, the ENET includes a two-dimensional exhaustive search of the tuning parameter pair, leading to the largest computation time. In comparison, the SLASSO simultaneously estimates the noise and coefficient, and its $\tt{R}$ package $\tt{scalreg}$ provides faster convergence than other penalized model selectors under consideration.
Perturbation Scheme II (Figure 2)
Figure 2: Perturbation Scheme II with (n,p)=(100,1000), $\kappa_{\text{Linear},\textbf{X}}$=10 and $\textbf{X}=\textbf{X}_{\text{Sc(0.8)}}$. Panel (A): power of detection; Panel (B): FPR of detection; Panels (C) and (D): MSE in prediction and probability that a selected submodel contains the true model before and after the removal of detected influential observations; Panel (E): algorithm running time in seconds.
@@
    Power: The detection power of the two existing methods, MIP and DF(LASSO) are no more than 0.10, considerably falling behind those of the newly proposed ClusMIP procedure across all simulation settings. While the poor performance of DF(LASSO) is dues to its leave-one-out design, for MIP, we postulate that contaminating the predictors while keeping the responses intact is unlikely to impose sufficient impact on the marginal correlation-based HIM measure, leading to lower detection power.
    When $\kappa_{\text{Linear},\textbf{X}}$ is moderate and for all types of design matrices, $\text{Power}_{\text{ClusMIP}(\text{LASSO})}$, $\text{Power}_{\text{ClusMIP}(\text{SLASSO})}$, $\text{Power}_{\text{ClusMIP}(\text{SCAD})}$ and $\text{Power}_{\text{ClusMIP}(\text{MCP})}$ perform best, and all approximately exceed 0.90. In contrast, $\text{Power}_{\text{ClusMIP}(\text{ENET})}$ is the worst performer in the ClusMIP family, showing larger variation over $\zeta$. As seen in Panel (A) of Figure 2, $\text{Power}_{\text{ClusMIP}(\text{ENET})}$ ranges between 0.65 and 0.85 in that specific simulation setting. One contributing factor is the need to select two tuning parameters for ENET may complicate and impede the formation of the GDF measure (3) to accurately exhibit the effects of influential observations. Indeed, poor performance of ClusMIP(ENET) in detection seems to be observed across all the simulation settings.
@@
    MSE of Prediction and Variable Selection: Similar to the conclusion from Perturbation Scheme I, the MSE of prediction and selection probability after removing outliers depends on the detection performance. Specifically, better detection results lead to lower MSE and higher selection probability after outlier removal. For example, as seen in Panels (C) and (D) of Figure 2, severe under-performance of DF(LASSO) in detection causes the highest MSE (between 8.96 and 9.42) and lowest selection probability (always 0). In contrast, ClusMIP(SCAD) and ClusMIP(SLASSO) attain the lowest MSE (between 0.73 and 0.94, and 1.29 and 1.73, respectively) and the highest selection probability (between 0.98 and 1.00, and 0.96 and 1.00, respectively) after removal of detected influential points.
@@
    FPR and Computation Time: they are similar to those under Perturbation Scheme I.
Perturbation Scheme III
The patterns of detection power, FPR, computation time, MSE of prediction and selection probability are analogous to the ones of Perturbation Scheme I.
@@ 6 Analysis of Neurologic Signature of Physical Pain
We apply existing procedures, as well as the newly proposed ClusMIP procedures to the fMRI data described in Section 2. We stress that even though the study [21] focuses on pain prediction, the pitfalls of ignoring outliers on downstream analysis have been thoroughly discussed both from scientific angles and via a synthetic example in Section 1. In particular, we have shown that the presence of multiple outliers significantly distorts variable selection and mitigates predictive power. Therefore, we focus on detecting influential points in Section 6.1. The benefits of excluding the identified outliers on variable selection, coefficient estimation, and prediction are then systematically examined in Sections 6.2 to 6.4. Based on our analyses, we conclude by providing concrete advice on the practical use of the newly proposed procedure for application to the thermal pain dataset [21] in Section 6.5.
Two metrics are considered to assess detection outcome: $\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}$ and $|\widehat{\text{I}}^{\text{infl}}_{\text{Detection}(\text{Selector})}\rvert$. They refer to the estimated index set of influential points and its cardinality, respectively. To obtain them, the existing methods, MIP and DF(LASSO), and the ClusMIP procedure applied to LASSO, SLASSO, ENET, SCAD and MCP are performed. Values for these two metrics for the different detection procedures are reported in Table 1. Since 10-fold CV is used for selecting tuning parameters and every choice is random, the analysis is repeated 1000 times. Since LASSO is used for both DF(LASSO) and ClusMIP(LASSO), to avoid ambiguity, results for both procedures are shown separately in Figures 4 and 5.
6.1 Detection of Influential Observations